# VISTA: Knowledge-Driven Interpretable Vessel Trajectory Imputation via Large Language Models

### v1

@@ 3.1. Data and Notation
An AIS record $x=(\iota,\lambda,\phi,\tau,\psi,\theta,s,\eta,\chi,\\ d,\ell,\beta,\kappa)$ consists of multiple attributes. At its core, an AIS record is anchored by the spatio-temporal attributes and vessel identifier $(\iota,\lambda,\phi,\tau)$, representing vessel identifier, longitude, latitude, and timestamp, which jointly specify the vessel’s geographic position at a particular time. Other attributes act as auxiliary descriptors:
@@
3.2. Underlying Knowledge
3.2.1. The definition of underlying knowledge
Definition 0 (Underlying Knowledge in Vessel Trajectory Imputation).
Given AIS data $\mathcal{X}$, the underlying knowledge $\mathcal{U}$ is defined as the integration of:
  • Structured Data-derived Knowledge (SDK) $\mathcal{K}_{d}$, distilled from AIS data $\mathcal{X}$, capturing vessel attributes, behavior patterns, imputation functions, and their empirical dependencies.
  • Implicit LLM Knowledge $\mathcal{K}_{\ell}$, consisting of maritime conventions and tacit operational know-how, together with commonsense knowledge embedded in LLM parameters.
The complete underlying knowledge is then given by
(2) $\mathcal{U}=\Phi(\mathcal{K}_{d},\mathcal{K}_{\ell}),$
where $\Phi$ denotes the process by which SDK $\mathcal{K}_{d}$ serves as trigger to activate and integrate relevant implicit knowledge from $\mathcal{K}_{\ell}$.
As shown in Figure 2, the construction of underlying knowledge $\mathcal{U}$ relies on two complementary sources. First, the SDK $\mathcal{K}_{d}$ is obtained from AIS data $\mathcal{X}$ through LLM-assisted analysis, where entities and their semantic relations can be distilled into a graph structure. Second, the implicit LLM knowledge $\mathcal{K}_{\ell}$ originates from the large-scale Internet corpus. Unlike $\mathcal{K}_{d}$, this knowledge cannot be explicitly enumerated as graph structure, but can be retrieved and aligned through queries triggered by $\mathcal{K}_{d}$. In this way, the structured component $\mathcal{K}_{d}$ serves as a trigger to activate and integrate relevant priors and commonsense information from $\mathcal{K}_{\ell}$, together yielding the complete underlying knowledge $\mathcal{U}$.
Figure 2. Illustration of the underlying knowledge.
3.2.2. The definition of structured data-derived knowledge graph
To enable trajectory imputation to be performed both efficiently and accurately, while also supporting downstream analytics, it is essential to carefully store and manage the SDK $\mathcal{K}_{d}$. Our design is guided by two considerations: 1) Human interpretability and LLM usability, ensuring that the knowledge is not only understandable to domain experts but also directly usable as triggers for LLMs; 2) Imputation accuracy and efficiency, binding vessel- and context-specific information to executable imputation functions with machine-actionable semantics (applicability conditions and parameters), while compressing massive AIS data into compact yet representative knowledge units for scalable retrieval and execution. To address these requirements, we introduce Structured Data-derived Knowledge Graph (SD-KG) to organize SDK.
@@ 3.2.2. The definition of structured data-derived knowledge graph
The SD-KG is represented as a graph $\mathcal{G}_{d}=(\mathcal{V}_{d},\mathcal{E}_{d})$, distilled from AIS data $\mathcal{X}$, where $\mathcal{V}_{d}$ denotes the set of nodes and $\mathcal{E}_{d}$ denotes their semantic relations. The node set $\mathcal{V}_{d}=\{\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f}\}$ is divided into:
  • Static Attribute Nodes $\mathcal{V}_{s}=\{\mathcal{V}_{\iota},\mathcal{V}_{\eta},\mathcal{V}_{\chi},\mathcal{V}_{d},\mathcal{V}_{\ell},\mathcal{V}_{\beta},\mathcal{V}_{\kappa},\mathcal{V}_{\sigma}\}$, mostly derived from the static and status-related attributes of AIS records (Definition 3.1). For each attribute type, a corresponding node set is constructed, where each node represents one possible value of that attribute. Specifically, $\mathcal{V}_{\sigma}$ differs from other sets in $\mathcal{V}_{s}$ as it is derived from core attributes and represents spatial context categories such as shipping lanes, ports, or anchorages.
  • Behavior Pattern Nodes $\mathcal{V}_{b}$ representing characteristic behavior patterns extracted from AIS data. Each node $v_{b}\in\mathcal{V}_{b}$ is defined as a tuple $v_{b}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$, where $p^{s}$ denotes the speed pattern, $p^{\theta}$ the course pattern, $p^{\psi}$ the heading pattern, $p^{i}$ the navigation intent, and $p^{\tau}$ the duration time. These patterns are primarily inferred from the kinematic and core attributes.
  • Imputation Function Nodes $\mathcal{V}_{f}$ representing available imputation methods. Each node $v_{f}\in\mathcal{V}_{f}$ is defined as a tuple $v_{f}=(f,d(f))$, where $f$ denotes an imputation function and $d(f)$ denotes its descriptive information. These nodes are primarily linked to the types of behavior patterns they can address, which are inferred from the core attributes of AIS records.
The edge set $\mathcal{E}_{d}=\{\mathcal{E}_{sb},\mathcal{E}_{bf}\}$ consists of two types of relations:
  • Static–Behavior edges $\mathcal{E}_{sb}\subseteq\mathcal{V}_{s}\times\mathcal{V}_{b}$: connecting static attribute nodes and behavior pattern nodes. For an edge $e=(v_{s},v_{b},\\ w)\in\mathcal{E}_{sb}$, the weight $w$ indicates how frequently the static attribute value $v_{s}$ co-occurs with the behavior pattern $v_{b}$ in $\mathcal{X}$.
  • Behavior–Function edges $\mathcal{E}_{bf}\subseteq\mathcal{V}_{b}\times\mathcal{V}_{f}$: connecting behavior pattern nodes and imputation function nodes. For an edge $e=(v_{b},v_{f},w)\in\mathcal{E}_{bf}$, the weight $w$ denotes the frequency with which the imputation method $v_{f}$ successfully imputes trajectories exhibiting the behavior pattern $v_{b}$.
Figure 3. Example of SD-KG.
@@
As shown in Figure 3, consider $\mathcal{G}_{d}=(\{\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f}\},\{\mathcal{E}_{sb},\\ \mathcal{E}_{bf}\})$ with static attribute nodes $\mathcal{V}_{s}=\{\mathcal{V}_{\kappa},\mathcal{V}_{\eta}\}$, behavior pattern nodes $\mathcal{V}_{b}=\{v_{b,1},v_{b,2},v_{b,3}\}$, and imputation function nodes $\mathcal{V}_{f}=\{v_{f,1},v_{f,2},v_{f,3}\}$. The pattern $v_{b,3}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$ corresponds to stable speed ($p^{s}$), gradual course change ($p^{\theta}$), low heading fluctuation ($p^{\psi}$), lane-following intent ($p^{i}$), and duration 1300–1350s ($p^{\tau}$). The edge $(v_{\eta,3},v_{b,3},6)\in\mathcal{E}_{sb}$ (red edge in Figure 3) shows that vessels with status “under way using engine” co-occurred with this pattern six times, while $(v_{b,3},v_{f,1},6)\in\mathcal{E}_{bf}$ (blue edge in Figure 3) indicates that the Linear Interpolation function $v_{f,1}=(f,d(f))$ successfully imputed this pattern six times.
This design guarantees the key requirements identified earlier.
Human interpretability and LLM usability. The construction achieves interpretability by explicitly organizing knowledge into semantically meaningful node types—static attributes, behavioral patterns, and imputation functions, that are readily understood by domain experts, while also providing deterministic structures that serve as precise triggers for $\mathcal{K}_{l}$. The decomposition of behaviors into speed, course, heading, intent, and duration enhances transparency, and the explicit pairing of imputation functions with descriptive semantics enables both humans and LLMs to reason about why a given function is appropriate for a specific pattern.
Imputation accuracy and efficiency. Accuracy is ensured in two complementary ways. First, the set of imputation functions $\mathcal{V}_{f}$ is derived directly from real AIS data, capturing empirically validated and executable strategies that can be invoked to reproduce vessel behaviors under similar conditions. Second, edge weights quantify the statistical association between attributes, patterns, and functions, thereby guiding the selection of imputation methods with data-grounded priors rather than unconstrained heuristics. Efficiency is achieved by compressing massive AIS data into a compact, executable knowledge graph representation, enabling inference through lightweight subgraph retrieval and direct function execution, while supporting incremental updates without retraining.
Together, these mechanisms ensure that $\mathcal{G}_{d}$ provides an accurate, efficient, and interpretable foundation for knowledge-driven vessel trajectory imputation. Further details on the construction and maintenance of SD-KG are provided in Section 4.2.
Figure 4. Partition of the AIS record sequence $\mathbf{X}_{\iota}$ into minimal segments (4 records per segment).
@@ 3.3. Problem Definition
We formalize the problem in three parts. With AIS records defined (Definitions 3.13.2), we first capture data incompleteness via minimal segments and an observation mask. Next, for complete AIS data, each segment is distilled into a knowledge unit and used to update the SD-KG $\mathcal{G}_{d}$. Finally, for incomplete segments, We define knowledge-driven trajectory imputation that queries the SD-KG $\mathcal{G}_{d}$ under contextual constraints to retrieve candidate behavior patterns and imputation functions; and then we selects and applies the most plausible function to reconstruct the segment, and returns the imputed segment together with justification composed of SD-KG statistics and grounded rationales.
@@ 3.3.1. Incomplete AIS Data
Trajectory missing in AIS data typically occurs in the form of block missing (Łubkowski and Laskowski, 2017), where consecutive records are absent rather than isolated points. To capture this phenomenon in a structured way, we introduce the concept of a minimal segment, which serves as the atomic block for modeling both observed and missing intervals.
@@
For a vessel-specific AIS record sequence $\mathbf{X}_{\iota}=\langle x_{1},x_{2},\dots,x_{T}\rangle$, a minimal segment $\mathbf{S}_{[t,t+m-1]}$ is defined as a consecutive block of AIS records
@@
where $m$ denotes the fixed segment length, corresponding to the minimal unit for modeling missing trajectory data (default $m=20$).
Thus, an AIS record sequence $\mathbf{X}_{\iota}=\langle x_{1},x_{2},\dots,x_{T}\rangle$ can be uniquely partitioned into an ordered concatenation of disjoint minimal segments
@@
where each $\mathbf{S}_{\iota}^{k}=\mathbf{S}_{[t_{k},\,t_{k}+m-1]}$ denotes the $k$-th minimal segment of fixed length $m$, ensuring that all segments are non-overlapping and together cover the entire AIS sequence of vessel $\iota$.
@@
As shown in Figure 4, let the fixed segment length be $m=4$. Around index $t$, the AIS record sequence $\mathbf{X}_{\iota}$ is partitioned into three consecutive minimal segments: $\mathbf{S}_{\iota}^{k-1}=\langle x_{t-4},x_{t-3},x_{t-2},x_{t-1}\rangle$, $\mathbf{S}_{\iota}^{k}=\langle x_{t},x_{t+1},x_{t+2},x_{t+3}\rangle$, and $\mathbf{S}_{\iota}^{k+1}=\langle x_{t+4},x_{t+5},x_{t+6},x_{t+7}\rangle$.
Definition 0 (Observation Mask).
Given the fixed-length partition $\mathbf{X}_{\iota}=\mathbf{S}_{\iota}^{1}\,\|\,\mathbf{S}_{\iota}^{2}\,\|\,\dots\,\|\,\mathbf{S}_{\iota}^{K}$, the segment mask is the binary vector
(5) $\mathbf{M}_{\iota}=\langle M_{1},\dots,M_{K}\rangle,$
where

### v2

@@ 3.1. Data and Notation
An AIS record $x=(\iota,\lambda,\phi,\tau,\psi,\theta,s,\eta,\chi,\\ d,\ell,\beta,\kappa)$ has multiple attributes. At its core, an AIS record is anchored by a vessel identifier and spatio-temporal attributes $(\iota,\lambda,\phi,\tau)$, representing vessel identifier, longitude, latitude, and timestamp, identifying a vessel and giving its geographic position at a particular time. Other attributes act as auxiliary descriptors:
@@
3.2. Structured Data-derived Knowledge
LLM outputs are prone to hallucination and fabrication (Shankar and others, 2024; Agrawal et al., 2024b). To produce trustworthy repair provenance, VISTA grounds LLM reasoning in Structured Data-derived Knowledge (SDK): a set of data-verifiable components—vessel attributes, behavior patterns, imputation functions, and their statistical associations—that are extracted from AIS data $\mathcal{X}$, validated through execution and statistical testing. As shown in Figure 2, this data-derived and continuously updated knowledge complements an LLM’s implicit knowledge acquired through pre-training on internet corpora and fixed at deployment. During imputation, validated SDK serves as grounded context that steers LLM reasoning toward faithful causal explanations, while the LLM contributes contextual reasoning and text generation capabilities.
Figure 2. Role of SDK in grounding LLM reasoning for trajectory imputation with repair provenance.
To enable efficient storage, querying, and retrieval of SDK, we introduce the Structured Data-derived Knowledge Graph (SD-KG).
@@ 3.2. Structured Data-derived Knowledge
An SD-KG is a graph $\mathcal{G}_{d}=(\mathcal{V}_{d},\mathcal{E}_{d})$, distilled from AIS data $\mathcal{X}$. The node set $\mathcal{V}_{d}=\mathcal{V}_{s}\cup\mathcal{V}_{b}\cup\mathcal{V}_{f}$ is composed as follows:
  • Static Attribute Nodes $\mathcal{V}_{s}=\mathcal{V}_{\iota}\cup\mathcal{V}_{\eta}\cup\mathcal{V}_{\chi}\cup\mathcal{V}_{d}\cup\mathcal{V}_{\ell}\cup\mathcal{V}_{\beta}\cup\mathcal{V}_{\kappa}\cup\mathcal{V}_{\sigma}$, mostly derived from the static and status-related attributes of AIS records (Definition 3.1). For each attribute, a corresponding node subset is constructed, where each node represents one possible value of that attribute. $\mathcal{V}_{\sigma}$ differs from other sets in $\mathcal{V}_{s}$ as it is derived from core attributes and represents spatial context categories such as shipping lanes, ports, or anchorages.
  • Behavior Pattern Nodes $\mathcal{V}_{b}$ representing characteristic behavior patterns extracted from AIS data. Each node $v_{b}\in\mathcal{V}_{b}$ is defined as a tuple $v_{b}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$, where $p^{s}$ is the speed pattern, $p^{\theta}$ the course pattern, $p^{\psi}$ the heading pattern, $p^{i}$ the navigation intent, and $p^{\tau}$ the duration. These patterns are primarily inferred from the kinematic and core attributes.
  • Imputation Function Nodes $\mathcal{V}_{f}$ representing available imputation methods. Each node $v_{f}\in\mathcal{V}_{f}$ is defined as a tuple $v_{f}=(f,d(f))$, where $f$ is an imputation function and $d(f)$ is its descriptive information. These nodes are primarily linked to the types of behavior patterns they can address, which are inferred from the core attributes of AIS records.
The edge set $\mathcal{E}_{d}=\mathcal{E}_{sb}\cup\mathcal{E}_{bf}$ consists of two types of edges:
  • Static–Behavior edges $\mathcal{E}_{sb}\subseteq\mathcal{V}_{s}\times\mathcal{V}_{b}$: connecting static attribute nodes and behavior pattern nodes. For an edge $e=(v_{s},v_{b},w)\\ \in\mathcal{E}_{sb}$, the weight $w$ indicates how frequently the static attribute value $v_{s}$ co-occurs with the behavior pattern $v_{b}$ in $\mathcal{X}$.
  • Behavior–Function edges $\mathcal{E}_{bf}\subseteq\mathcal{V}_{b}\times\mathcal{V}_{f}$: connecting behavior pattern nodes and imputation function nodes. For an edge $e=(v_{b},v_{f},w)\in\mathcal{E}_{bf}$, the weight $w$ is the frequency with which the imputation method $v_{f}$ successfully imputes trajectories exhibiting the behavior pattern $v_{b}$.
Figure 3. Example of SD-KG $\mathcal{G}_{d}$.
@@
As shown in Figure 3, consider $\mathcal{G}_{d}=(\mathcal{V}_{s}\cup\mathcal{V}_{b}\cup\mathcal{V}_{f},\;\mathcal{E}_{sb}\cup\mathcal{E}_{bf})$ with static attribute nodes $\mathcal{V}_{s}=\mathcal{V}_{\kappa}\cup\mathcal{V}_{\eta}$, behavior pattern nodes $\mathcal{V}_{b}=\{v_{b,1},v_{b,2},v_{b,3}\}$, and imputation function nodes $\mathcal{V}_{f}=\{v_{f,1},v_{f,2},v_{f,3}\}$. The pattern $v_{b,3}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$ corresponds to stable speed ($p^{s}$), gradual course change ($p^{\theta}$), low heading fluctuation ($p^{\psi}$), lane-following intent ($p^{i}$), and duration 1300–1350s ($p^{\tau}$). The edge $(v_{\eta,3},v_{b,3},6)\in\mathcal{E}_{sb}$ (red edge in Figure 3) shows that vessels with status “under way using engine” co-occurred with this pattern six times, while $(v_{b,3},v_{f,1},6)\in\mathcal{E}_{bf}$ (blue edge in Figure 3) indicates that the Linear Interpolation function $v_{f,1}=(f,d(f))$ successfully imputed this pattern six times.
The SD-KG design is driven by two goals: making the knowledge both human-interpretable and LLM-usable, and ensuring that it supports accurate and efficient imputation.
Human interpretability and LLM usability. Semantically meaningful node types—static attributes, behavior patterns, and imputation functions—are readily understood by domain experts and simultaneously provide structured context that grounds LLM reasoning. The decomposition of behaviors into speed, course, heading, intent, and duration enhances transparency, and pairing imputation functions with descriptive semantics enables both humans and LLMs to reason about function–pattern appropriateness.
Imputation accuracy and efficiency. Imputation functions in $\mathcal{V}_{f}$ are derived from real AIS data, capturing empirically validated and executable strategies, while edge weights quantify statistical associations to guide function selection with data-grounded priors. Efficiency is achieved by compressing massive AIS data into a compact graph, enabling lightweight subgraph retrieval and direct function execution with incremental updates. Further details on the construction and maintenance of SD-KG are provided in Section 4.2.
Figure 4. Partition of an AIS record sequence $\mathbf{X}_{\iota}$ into minimal segments (4 records per segment).
@@ 3.3. Problem Definition
We formalize the problem in three parts. With AIS records defined (Definitions 3.13.2), we first capture data incompleteness via minimal segments and an observation mask. Next, for complete AIS data, each segment is distilled into a knowledge unit and used to update the SD-KG $\mathcal{G}_{d}$. Finally, for incomplete segments, we define knowledge-driven trajectory imputation that queries $\mathcal{G}_{d}$ under contextual constraints to retrieve candidate behavior patterns and imputation functions, selects and applies the most plausible function to reconstruct the segment, and returns the imputed segment together with its repair provenance.
@@ 3.3.1. Incomplete AIS Data
Missing data in AIS trajectories typically occurs as contiguous block-level gaps (Łubkowski and Laskowski, 2017), where consecutive records are absent rather than isolated points. To capture this phenomenon in a structured way, we introduce the concept of a minimal segment, which serves as the atomic block for modeling both observed and missing intervals.
@@
For a vessel-specific AIS record sequence $\mathbf{X}_{\iota}=\langle x_{1},x_{2},\dots,x_{T}\rangle$, a minimal segment $\mathbf{S}_{[t,t+m-1]}$ is defined as a consecutive block of AIS records:
@@
where $m$ is the fixed segment length, corresponding to the minimal unit for modeling missing trajectory data (default $m=20$).
An AIS record sequence $\mathbf{X}_{\iota}=\langle x_{1},\dots,x_{T}\rangle$ can be partitioned uniquely into a concatenation of disjoint minimal segments:
@@
where each $\mathbf{S}_{\iota}^{k}=\mathbf{S}_{[t_{k},\,t_{k}+m-1]}$ is the $k$-th minimal segment of fixed length $m$, ensuring that all segments are non-overlapping and together cover the entire AIS sequence of vessel $\iota$.
Example 0.
Consider Figure 4, where the fixed segment length is $m=4$. Around index $t$, the AIS record sequence $\mathbf{X}_{\iota}$ is partitioned into three consecutive minimal segments: $\mathbf{S}_{\iota}^{k-1}=\langle x_{t-4},x_{t-3},x_{t-2},x_{t-1}\rangle$, $\mathbf{S}_{\iota}^{k}=\langle x_{t},x_{t+1},x_{t+2},x_{t+3}\rangle$, and $\mathbf{S}_{\iota}^{k+1}=\langle x_{t+4},x_{t+5},x_{t+6},x_{t+7}\rangle$.
Definition 0 (Observation Mask).
Given the fixed-length partition $\mathbf{X}_{\iota}=\mathbf{S}_{\iota}^{1}\,\|\,\mathbf{S}_{\iota}^{2}\,\|\,\dots\,\|\,\mathbf{S}_{\iota}^{K}$, the observation mask is the binary sequence
(4) $\mathbf{M}_{\iota}=\langle M_{1},\dots,M_{K}\rangle,$
where
(5) $M_{k}=\begin{cases}1&\text{if }\forall\,j\in[t_{k},t_{k}+m-1],~x_{j}\text{ has all attributes present}\\ 0&\text{otherwise}\end{cases}$
Example 0.
Continuing Example 3.6, using the convention $M_{k}=0$ for a segment with missing records and $M_{k}=1$ for a fully observed segment, the configuration in Figure 4 yields $M_{k-1}=1$, $M_{k}=0$, and $M_{k+1}=0$, i.e., the mask sequence $\mathbf{M}_{\iota}$ is $\langle\ldots,1,0,0,\ldots\rangle$.
3.3.2. Updating of Structured Data-derived Knowledge
A minimal segment $\mathbf{S}_{\iota}^{k}$ is a time-indexed sample and cannot directly update the SD-KG $\mathcal{G}_{d}$, which is organized over semantic node types and their relations. To bridge segment-level data and the graph-level representation, we introduce the knowledge unit as the minimal semantic unit distilled from a segment: it includes a static attribute, a behavior pattern, and an imputation function, and it serves as the atomic granularity for SD-KG updating.
Definition 0 (Knowledge Unit).
A knowledge unit $u$ is defined as a triple $u\;=(v_{s},v_{b},v_{f})\;=((\iota,\eta,\chi,d,\ell,\beta,\kappa,\sigma),\;(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau}),\;\\ (f,d(f)))$, where $v_{s}$ is a static-attribute vector, $v_{b}$ is a behavior pattern, and $v_{f}$ is an imputation method.
Given a minimal segment $\mathbf{S}_{\iota}^{k}$ with $M_{\iota}^{k}=1$, the updating of SD-KG refers to the following procedure:
  • Knowledge distillation: $\mathbf{S}_{\iota}^{k}\mapsto u_{\iota}^{k}=(v_{s},v_{b},v_{f})$, which compresses a raw segment into a static attribute $v_{s}$, a behavior pattern $v_{b}$, and an imputation method $v_{f}$.
  • Graph integration and maintenance: The node set $\mathcal{V}_{d}$ and edge set $\mathcal{E}_{d}$ of $\mathcal{G}_{d}$ are updated by inserting or merging $v_{s}$, $v_{b}$, and $v_{f}$ and by adjusting the corresponding edges $(v_{s},v_{b})\in\mathcal{E}_{sb}$ and $(v_{b},v_{f})\in\mathcal{E}_{bf}$. Redundant nodes are merged, and edge weights are aggregated to ensure that $\mathcal{G}_{d}$ remains compact and efficient for retrieval.
@@ 3.3.2. Updating of Structured Data-derived Knowledge
In Figure 5, there is a segment $\mathbf{S}_{\iota}^{k}=\langle x_{t},\ldots,x_{t+m-1}\rangle$ with $M_{\iota}^{k}=1$. Within this segment, the speed $s$ remains stable at approximately $5.4$ kn, the course $\theta$ varies gradually around $110^{\circ}$, the heading $\psi$ exhibits mild fluctuations near $110^{\circ}$, and both the navigation status and vessel type are constant, recorded as “under way using engine” and “cargo”.
Step 1: Knowledge distillation. From the static and status-related attributes, together with the vessel’s geographic position, we derive the static attributes $v_{s}=(\iota,\eta,\kappa,\sigma)$ (listing only the attributes relevant to this example for brevity), describing a cargo vessel $\iota$ operating “under way using engine” near the Traffic Separation Scheme in the Great Belt East Channel. From the kinematic attributes, we extract the behavior pattern $v_{b}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$, corresponding to stable speed, gradual course change, mild heading fluctuation, lane-following intent, and duration 1300130013501350s. To avoid redundancy in $\mathcal{V}_{b}$, the states $p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau}$ are preferentially aligned with existing entries, in this case matching $v_{b,3}\in\mathcal{V}_{b}$ (continuing Example 3.4). Finally, based on the trajectory trend, we generate the imputation method $v_{f}=(f,d(f))$, where $f$ is the executable function that simulates the observed dynamics and $d(f)$ records its description.
Step 2: Graph integration and maintenance. Continuing Example 3.4, $v_{b}$ coincides with $v_{b,3}$, and the imputation function $v_{f}$ is semantically equivalent to $v_{f,1}$. The graph is updated accordingly: the weights of edges $(v_{\eta,3},v_{b,3})$ and $(v_{b,3},v_{f,1})$ are each incremented by $1$ (from $6$ to $7$, red edges), while a new edge $(v_{\kappa,2},v_{b,3})$ is added with weight $1$ (blue edge).
Further details on SD-KG updates are presented in Section 4.2.
3.3.3. Knowledge-Driven Trajectory Imputation
Before specifying the imputation procedure, we formalize the target output. The outcome should not only reconstruct the missing trajectory but also produce structured repair provenance that documents the reasoning behind each repair decision.
Definition 0 (Imputation Outcome with Repair Provenance).
For a minimal segment $\mathbf{S}_{\iota}^{k}$ with $M_{\iota}^{k}=0$, the imputation outcome is defined as $R_{\iota}^{k}=(\widehat{\mathbf{S}}_{\iota}^{k},\;\mathcal{J}_{\iota}^{k})$, where $\widehat{\mathbf{S}}_{\iota}^{k}$ is the imputed trajectory segment and $\mathcal{J}_{\iota}^{k}$ is the accompanying repair provenance. The repair provenance $\mathcal{J}_{\iota}^{k}$ consists of:
  • What — behavior-pattern estimation and rationale $(v_{b}^{*},\,\mathcal{J}_{\iota}^{k,b})$: $v_{b}^{*}$ is the estimated behavior pattern, and $\mathcal{J}_{\iota}^{k,b}$ provides contextual and statistical evidence from $\mathcal{G}_{d}$ supporting this estimation.
  • How — function selection and rationale $(v_{f}^{*},\,\mathcal{J}_{\iota}^{k,f})$: $v_{f}^{*}$ is the imputation function applied to obtain $\widehat{\mathbf{S}}_{\iota}^{k}$, and $\mathcal{J}_{\iota}^{k,f}$ contains statistical evidence from $\mathcal{G}_{d}$ justifying the choice of this function.
  • Why — causal explanation $\mathcal{J}_{\iota}^{k,h}$: a natural-language explanation of why the vessel exhibits the estimated behavior, grounded in the SDK context and generated by an LLM.
Example 0.